OpenStudy (rainbow_dashie):

Fan, medal, and cookies Which equations show that the set of whole numbers is not closed under subtraction?

OpenStudy (anonymous):

is there more too this question?

OpenStudy (rainbow_dashie):

Choose all answers that are correct. A. 1 – (–2) = 3 B. 1 – 2 = –1 C. 2 – 0 = 2 D. 2 – 4 = –2

OpenStudy (rainbow_dashie):

A. 1-(-2)=3

OpenStudy (rainbow_dashie):

B.1-2=-1

OpenStudy (rainbow_dashie):

C.2-0=2

OpenStudy (rainbow_dashie):

D.2-4=-2

OpenStudy (rainbow_dashie):

Choose all answers that are correct

OpenStudy (rainbow_dashie):

Nobody wants cookies

OpenStudy (anonymous):

c and d? maybe dont trust me....i suck at math!

OpenStudy (rainbow_dashie):

@uChezzy @cjackson2002 @Walkerman15 @aliboo997 @kyley_bear

OpenStudy (anonymous):

I have no clue hun let me look on google:)

OpenStudy (rainbow_dashie):

Ok dont call me hun:)

OpenStudy (anonymous):

dont call my sister hun!! got that! :@

OpenStudy (anonymous):

I call everyone hun lol what is wrong with that.. lol I don't have to go look for you

OpenStudy (rainbow_dashie):

I was messing my sis is mean

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

aw cmon take a joke and im sorry it didnt sound or look like that

OpenStudy (rainbow_dashie):

stop being mean sis

OpenStudy (anonymous):

im noooot :(

OpenStudy (anonymous):

im a nice do gooder

OpenStudy (anonymous):

it is okay lol I forgive you lol it didn't look like a joke but if you say so I can trust you both:)

OpenStudy (anonymous):

zooglobble

OpenStudy (anonymous):

:P lol

OpenStudy (anonymous):

lol huh?

OpenStudy (anonymous):

eyupies :)

OpenStudy (rainbow_dashie):

Dont listen to my sister

OpenStudy (anonymous):

lol okay ... well I could not find an answer and I don't understand this question either why don't you go to the members online tab and tag the guys/girls at the top:) they should be able to help you :)

OpenStudy (anonymous):

hey!!!!!!!! :(:(:(:(

OpenStudy (rainbow_dashie):

Im smarter

OpenStudy (rainbow_dashie):

Thx for trying @kyley_bear

OpenStudy (anonymous):

yuppers :)

OpenStudy (anonymous):

stop being mean. now what was your question kyley_bear?

OpenStudy (anonymous):

o wait u dnt have question srry:)

OpenStudy (rainbow_dashie):

@mathmath333 @marylou004

OpenStudy (anonymous):

hu r they?

OpenStudy (rainbow_dashie):

@leonardo0430

OpenStudy (leonardo0430):

too much chat i cant even find the question XD

OpenStudy (anonymous):

haha

OpenStudy (leonardo0430):

the answer is D. did it real quick

OpenStudy (rainbow_dashie):

ok thx

OpenStudy (anonymous):

teeheehee

OpenStudy (anonymous):

https://www.youtube.com/watch?v=v4RGhaL8dcs must watch! so funny!

OpenStudy (anonymous):

comment on it!